# Distributed Quantum Approximate Optimization Algorithm on Integrated High-Performance Computing and Quantum Computing Systems for Large-Scale Optimization

### v1

@@ 2.2 Performance Analysis of DQAOA
Figure 3: Performance analysis of DQAOA. Accuracy and time-to-solution for a 150-bit problem as a function of (a) the number of iterations, and (b) the number of cores used. (c) Accuracy and (d) time-to-solution of dq-QAOA and DQAOA as a function of the problem size (n). Note that the number of iterations, sub-QUBO size, and the number of cores used for DQAOA are set to 30, 4, and n, respectively.
@@ 2.3 Active Learning with DQAOA for Real-World Optimization Problems
We further develop an active learning algorithm working with our DQAOA (AL-DQAOA) for optimizing material structures (layered metamaterials for spectral filters), featuring complex and large optimization spaces. This AL-DQAOA integrates machine learning, our DQAOA, and wave-optics simulation in an iterative loop (Fig. 4a). Here, a machine learning model (factorization machine, FM) is used to generate a QUBO as a surrogate, describing the relationship between material structures and their figure-of-merits (FOMs) [36, 37]. Then, DQAOA solves the given QUBO problem to predict an optimal state that represents an optimal material structure ($\textit{x}_{po}$). Subsequently, optical properties are evaluated by wave-optics simulation to calculate an FOM of the identified material structure ($\textit{FOM}_{po}$). Finally, the dataset previously used for FM training is updated by including $\textit{x}_{po}$ and $\textit{FOM}_{po}$. Iterations of these processes (noted as optimization cycles) improve the dataset quality and refine QUBO surrogates, gradually converging toward the optimal state in the global optimization space. It should be noted that this algorithm is designed for minimization optimization problems, where materials with lower FOMs yield higher performances (see the Method section for more details).
@@
We apply AL-DQAOA to optimize spectral filters with complex optical properties depending on the wavelength of the incident light, which can be used as a transparent radiative cooler (TRC) for window applications (Figs. 4b and 4c, see the Method section for more details). AL-DQAOA can effectively find the optimal structure in a global optimization space for a 6-layered structure (12-bit optimization problem), as shown in Figs. 4d and S6. More iterations are required for larger optimization spaces, generally achieving convergence within several thousands of optimization cycles (Figs. 4d and S7). Here, the optimization time per cycle for 12, 30, 50, and 100-bit systems is 13, 20, 29, and 35 s, respectively, with the simulation time using the transfer matrix method (TMM) being less than 0.1 s (Fig. 4e) [38]. Note that DQAOA takes 7 to 10 s and the total time required for completing optimization of 12, 30, 50, and 100-bit systems are respectively $\sim$0.9, $\sim$11.1, $\sim$32.9, and $\sim$48.9 hours, which should be much faster than AL with dq-QAOA (AL-dq-QAOA). To further demonstrate the superior performance of DQAOA, we optimize the same systems using dq-QAOA instead of DQAOA within AL. The results show that optimization using AL-dq-QAOA results in higher FOMs (indicating lower performances) and longer optimization times since dq-QAOA has lower accuracy and takes longer in solving given QUBOs than DQAOA (Figs. S8 and S9), which agree well with our performance analysis study in Figs. 2 and 3. These results highlight the benefit of using DQAOA within the AL algorithm not just for reducing optimization time but also for identifying high-performance materials.
@@
Figure 4: The AL-DQAOA algorithm for solving large-scale optimization problems in material science, exemplified by spectral filters with 12, 30, 50, and 100-bit systems. a. The schematic of AL-DQAOA integrating machine learning, DQAOA, and performance evaluation in an iterative loop. b. The schematic of a spectral filter for TRC windows, where the planar multilayered structure (PML) is subject to optimization. c. Solar spectral irradiance (yellow shade, air mass 1.5 global) and transmitted irradiance through the ideal spectral filter (red shade). The ideal spectral filter transmits the solar spectrum only in the visible range. d. The evolution of FOM as a function of optimization cycles for different problem sizes. e. Time required for an optimization cycle of AL-DQAOA for different problem sizes. f. Calculated optical properties of the spectral filter optimized by AL-DQAOA. The yellow shade presents solar spectral irradiance. g. Energy-saving potential when using the optimized spectral filter as TRC windows in cities around the world, compared to a conventional window.
@@ 3.   Conclusion
In this work, we have developed DQAOA leveraging the integrated HPC-QC system to solve large-scale, real-world optimization problems based on conventional quantum algorithms and conventional quantum computing systems. Our performance analysis demonstrates that DQAOA achieves high accuracy and significantly reduces time-to-solution by employing distributed computing across multi-cores and nodes on the HPC-QC integrated system. Consequently, DQAOA can effectively handle extensive problems, such as a 1,000-bit problem, achieving an accuracy of $\sim$99and time-to-solution of $\sim$276 s, which conventional methods may find challenging to solve. Furthermore, we have proposed AL-DQAOA, an optimization algorithm aiming to optimally design material structures leveraging gate-based quantum computing. This algorithm successfully optimizes photonic structures (spectral filters for TRC as examples), which represent 100-bit real-world optimization problems. Importantly, AL-DQAOA performs much more accurately and faster in identifying optimal structures compared to AL-dq-QAOA. This suggests utilizing DQAOA within AL for material design tasks if HPC-QC integrated systems are available. We present the optical properties of the optimized functional material, which resembles the ideal one, and demonstrate the energy-saving potential of the optimized material. We expect that the proposed DQAOA working with HPC-QC integrated systems can be generalized to various optimization tasks, such as finance, routing, scheduling, and assignment. In addition, the proposed AL-DQAOA will be useful for material design in general. Future enhancements could include adopting advanced quantum algorithms and computing systems to further optimize DQAOA’s performance and expand AL-DQAOA’s applicability in material design and beyond.
@@ 4.1 QUBO - Representing Optimization Problems
$\bar{\textbf{{x}}}={\rm arg\leavevmode\nobreak\ \min_{\textit{x}}}\leavevmode\nobreak\ \bar{y}$ (1)
$\bar{y}=\sum_{\textbf{{x}}_{i}\in\{0,1\}^{n}}\textbf{{x}}^{T}\mathbf{Q}\textbf{{x}}$ (2)
@@ 4.2 Building QUBOs
$y=w_{0}+\sum_{i=1}^{n}w_{i}{\textit{x}}_{i}+\frac{1}{2}\sum_{f=1}^{m}\left[\left(\sum_{i=1}^{n}v_{i,f}{\textit{x}}_{i}\right)^{2}-\sum_{i=1}^{n}v_{i,f}^{2}{\textit{x}}_{i}^{2}\right]$ (3)
@@
To analyze the performance of QAOA and DQAOA, we generate QUBO problems after training FM with datasets that depict the relationship between material structures (spectral filters) and their calculated FOMs. Hence, these QUBO matrices are large and dense (Fig. S5), reflecting real-world optimization scenarios. In this study, the QUBO size (i.e., the length of a binary bit string representing a material structure) ranges from 6 to 1,000, indicating 6-bit to 1,000-bit problems.
@@ 4.3 QUBO decomposition
Large QUBO problems are hard to solve with current conventional quantum technologies. Therefore, we decompose a large QUBO into smaller sub-QUBOs and aggregate sub-solutions from the sub-QUBOs to obtain a global solution. This decomposition strategy greatly reduces the complexity of handling a large QUBO. For example, a QUBO problem with 8 decision variables involves 8 self-interactions and 28 pairwise interactions. By fixing 3 decisions and decomposing the QUBO into several sub-QUBOs with 5 decisions each, we can reduce the complexity of each sub-QUBO to have 5 self-interactions and 10 interactions (Fig. 1c), making it more suitable for current quantum computing systems. This QUBO decomposition can find a true global optimal solution of an original large QUBO when preserving essential structure and interactions of the original problem [43, 44]. However, it can be computationally expensive and ineffective to solve all sub-QUBOs to identify the global optimal solution, especially for large-scale problems due to the exponential number of possible combinations ${}_{\textbf{{n}}}C_{\textbf{{k}}}$. Here, n and k represent the size of the original QUBO and sub-QUBOs, respectively. Hence, our algorithm iterates the decomposition and aggregate processes with randomly selected sub-QUBOs among the whole possibilities. Then, sub-QUBOs are solved using DQAOA.
4.4 DQAOA
@@
DQAOA can operate on both single-core and multi-cores/nodes within HPC-QC integrated systems. Our algorithm is tested on the Frontier supercomputer at the Oak Ridge Leadership Computing Facility, which is equipped with 64-core AMD “Optimized 3rd Gen EPYC” CPUs and 512 GB of memory per compute node [45]. In addition, each core interfaces with quantum systems (specifically, Qiskit Aer simulator in this study). dq-QAOA runs with a single core while DQAOA exploits multiple cores and nodes depending on the problem size. In this work, we use 50 cores per node, meaning that 150-bit problems would require 150 cores across 3 compute nodes. Distributed tasks across multi-cores and nodes are achieved by a message passing interface (MPI) implementation using mpi4py, enabling the simultaneous handling of numerous sub-QUBOs in each iteration [46].

### v2

@@ 2.2 Performance Analysis of DQAOA
Figure 3: Performance analysis of DQAOA. Approximation ratio and time-to-solution for a 150-bit problem as a function of (a) the number of iterations, and (b) the number of cores used. (c) Approximation ratio and (d) time-to-solution of dq-QAOA and DQAOA as a function of the problem size (n). Note that the number of iterations, sub-QUBO size, and the number of cores used for DQAOA are set to 30, 4, and n, respectively.
2.3 Performance Comparison and Implementation on the QCSC Architecture
While the divide-and-conquer approach with QAOA (DC-QAOA) has been proposed to address large-scale QUBO problems [27], it is less effective for dense problems as its policies do not consider many correlated interactions that exist between sub-problems. In contrast, the decomposition and aggregation policies introduced in DQAOA effectively address many correlated interactions. Fig. 4a shows that DQAOA achieves the best approximation ratio compared to that of QAOA, DC-QAOA, and dq-QAOA. Moreover, DQAOA significantly outperforms these methods in terms of time-to-solution, with orders of magnitude faster performance (Fig. 4b). The QUBOs used in this study represent real-world problems that are both large and dense (Fig. S5), and DQAOA uniquely demonstrates the capability to solve such problems with high efficiency. In particular, DQAOA’s flexibility may allow it to extend to higher-order unconstrained binary optimization (HOBO) problems using the same decomposition and aggregation policies. In contrast, other variable-fixing schemes, such as QBSolv provided by D-Wave Systems [36], may be limited to QUBO problems and lack support for distributed computing across multi-cores/nodes. Exploring the application of DQAOA to large and dense HOBO problems is a promising direction for future research, and will be addressed in our future work.
DQAOA has been successfully implemented on real quantum devices integrated with HPC systems (i.e., the QCSC architecture), demonstrating the feasibility of current quantum hardware for practical applications (Fig. 4c, see details in the Method section). Hardware noise and the extensive number of two-qubit gates required for solving a QUBO with n = 22 negatively impact algorithm performance [21]. Additionally, the sparse connectivity inherent to superconducting-based quantum devices further degrades performance [20], particularly for highly entangled problems common in real-world QUBO applications. Consequently, standard QAOA achieves a low approximation ratio of 0.6740 on the IBM-Strasbourg quantum device. This low performance is primarily due to hardware imperfections, as evidenced by relatively higher approximation ratio on noiseless simulators (Fig. S1a). However, despite the constrained number of iterations posed by limited quantum device access, DQAOA and dq-QAOA achieve significantly higher approximation ratios of 0.9727 and 0.9365, respectively, on the QCSC architecture (Fig. 4d). Notably, DQAOA yields the highest approximation ratio while requiring a shorter time-to-solution than dq-QAOA, emphasizing its superior performance. These results align well with those observed on emulators (FakeBrisbane and FakeKyiv; Fig. 4e).
To further evaluate scalability, we extend the problem size to a 100-bit QUBO, which is intractable for standard QAOA. Figure 4f shows that DQAOA and dq-QAOA can successfully solve this problem on the QCSC architecture. Both methods achieve high approximation ratios despite the restricted number of iterations, with DQAOA providing superior performance as expected (Fig. 4f). These results demonstrate DQAOA’s ability to efficiently tackle large and dense real-world optimization problems that conventional QAOA cannot handle, by leveraging the QCSC architecture and employing efficient decomposition and aggregation strategies. Furthermore, the performance of DQAOA is expected to further improve with the implementation of advanced (optimized) QAOA and next-generation quantum systems, which are beyond the scope of this study.
Figure 4: Performance analysis of dq-QAOA and DQAOA, and hardware implementation. Comparison of (a) approximation ratio and (b) time-to-solution between QAOA, DC-QAOA, dq-QAOA, and DQAOA, with DQAOA serving as the baseline for time-to-solution. c, Schematic representation of the implementation of dq-QAOA and DQAOA on the QCSC architecture. For a QUBO with n = 22, (d) approximation ratio and (e) time-to-solution of QAOA, dq-QAOA, and DQAOA on quantum devices (orange bar; IBM-Strasbourg and IBM-Kyiv) and emulators (grey bar; FakeBrisbane and FakeKyiv). f, For a QUBO with n = 100, approximation ratio of DQAOA (solid lines) and dq-QAOA (dotted lines) on quantum devices (orange; IBM-Strasbourg and IBM-Kyiv) and emulators (grey; FakeBrisbane and FakeKyiv).
@@ 2.4 Active Learning with DQAOA for Real-World Optimization Problems
We further develop an active learning algorithm working with our DQAOA (AL-DQAOA) for optimizing material structures (layered metamaterials for spectral filters), featuring complex and large optimization spaces. This AL-DQAOA integrates machine learning, our DQAOA, and wave-optics simulation in an iterative loop (Fig. 5a) within the QCSC architecture. Here, a machine learning model (factorization machine, FM) is used to generate a QUBO as a surrogate, describing the relationship between material structures and their figure-of-merits (FOMs) [37, 38]. Then, DQAOA solves the given QUBO problem to predict an optimal state that represents an optimal material structure ($\textit{x}_{po}$). Subsequently, optical properties are evaluated by wave-optics simulation to calculate an FOM of the identified material structure ($\textit{FOM}_{po}$). Finally, the dataset previously used for FM training is updated by including $\textit{x}_{po}$ and $\textit{FOM}_{po}$. Iterations of these processes (noted as optimization cycles) improve the dataset quality and refine QUBO surrogates, gradually converging toward the optimal state in the global optimization space. It should be noted that this algorithm is designed for minimization optimization problems, where materials with lower FOMs yield higher performances (see the Method section for more details).
@@
We apply AL-DQAOA to optimize spectral filters with complex optical properties depending on the wavelength of the incident light, which can be used as a transparent radiative cooler (TRC) for window applications (Figs. 5b and 5c, see the Method section for more details). AL-DQAOA can effectively find the optimal structure in a global optimization space for a 6-layered structure (12-bit optimization problem), as shown in Figs. 5d and S6. More iterations are required for larger optimization spaces, generally achieving convergence within several thousands of optimization cycles (Figs. 5d and S7). Here, the optimization time per cycle for 12, 30, 50, and 100-bit systems is 13, 20, 29, and 35 s, respectively, with the simulation time using the transfer matrix method (TMM) being less than 0.1 s (Fig. 5e) [39]. Note that DQAOA takes 7 to 10 s and the total time required for completing optimization of 12, 30, 50, and 100-bit systems are respectively $\sim$0.9, $\sim$11.1, $\sim$32.9, and $\sim$48.9 hours, which should be much faster than AL with dq-QAOA (AL-dq-QAOA). To further demonstrate the superior performance of DQAOA, we optimize the same systems using dq-QAOA instead of DQAOA within AL. The results show that optimization using AL-dq-QAOA results in higher FOMs (indicating lower performances) and longer optimization times since dq-QAOA has a lower approximation ratio and takes longer in solving given QUBOs than DQAOA (Figs. S8 and S9), which agree well with our performance analysis study in Figs. 2 and 3. These results highlight the benefit of using DQAOA within the AL algorithm not just for reducing optimization time but also for identifying high-performance materials.
@@
Figure 5: The AL-DQAOA algorithm for solving large-scale optimization problems in material science, exemplified by spectral filters with 12, 30, 50, and 100-bit systems. a, The schematic of AL-DQAOA integrating machine learning, DQAOA, and performance evaluation in an iterative loop. b, The schematic of a spectral filter for TRC windows, where the planar multilayered structure (PML) is subject to optimization. c, Solar spectral irradiance (yellow shade, air mass 1.5 global) and transmitted irradiance through the ideal spectral filter (red shade). The ideal spectral filter transmits the solar spectrum only in the visible range. d, The evolution of FOM as a function of optimization cycles for different problem sizes. e, Time required for an optimization cycle of AL-DQAOA for different problem sizes. f, Calculated optical properties of the spectral filter optimized by AL-DQAOA. The yellow shade presents solar spectral irradiance. g, Energy-saving potential when using the optimized spectral filter as TRC windows in cities around the world, compared to a conventional window.
@@ 3. Conclusion
In this work, we have developed DQAOA to solve large-scale, real-world optimization problems based on conventional quantum algorithms and current quantum computing systems. Our performance analysis demonstrates that DQAOA achieves a high approximation ratio and significantly reduces time-to-solution by employing distributed computing across multi-cores and nodes on the HPC-QC integrated system. Consequently, DQAOA effectively handles extensive problems, such as a 1,000-bit problem, achieving an approximation ratio of $\sim$99and time-to-solution of $\sim$276 s, which conventional methods may find challenging to solve. Furthermore, we demonstrate the implementation of DQAOA on the QCSC architecture, achieving the high-quality solution. To extend its application in material science, we have proposed AL-DQAOA, an advanced optimization algorithm aiming to optimally design material structures using gate-based QC. This algorithm successfully optimizes photonic structures (spectral filters for TRC as examples), representing 100-bit real-world optimization problems. We expect that the proposed DQAOA working with the QCSC architecture can be generalized to various optimization tasks, such as finance, routing, scheduling, and assignment. In addition, the proposed AL-DQAOA will be useful for material design in general. Future enhancements could include adopting advanced quantum algorithms and next-generation quantum computing systems, including simulators and devices, to further optimize DQAOA’s performance and expand AL-DQAOA’s applicability in material design and beyond.
@@ 4. Methods
4.1 QAOA
QAOA is a hybrid quantum-classical algorithm designed to solve combinatorial optimization problems by leveraging quantum systems. For minimization tasks, the objective is to prepare a quantum state that approximates the lowest energy state of a cost Hamiltonian ($H_{C}$), corresponding to the optimal solution of the classical problem. To achieve this, the classical problem is encoded into $H_{C}$, where binary decision variables ($d_{i}\in\{-1,1\}$) are mapped to the eigenvalues of Pauli-$Z$ operators ($\sigma^{z}$). The algorithm starts with an initial quantum state ($|\psi_{0}\rangle=|+\rangle^{\otimes n}$) as a uniform superposition using Hadamard gates applied to all qubits. QAOA then alternates between applying two operators: one derived from the problem Hamiltonian ($H_{C}$) and the other from a mixer Hamiltonian ($H_{M}=\sum_{i}\sigma^{x}_{i}$), which is a parameterized quantum state [17, 44]:
${|\psi(\beta,\gamma)\rangle=e^{-i\beta_{l}H_{M}}e^{-i\gamma_{l}H_{C}}\cdots e^{-i\beta_{1}H_{M}}e^{-i\gamma_{1}H_{C}}|\psi_{0}\rangle,}$ (1)
where $\beta=(\beta_{1},\beta_{2},\dots,\beta_{l})$ and $\gamma=(\gamma_{1},\gamma_{2},\dots,\gamma_{l})$ are variational parameters, and $l$ is the number of layers. These parameters are optimized classically to minimize the expected value of the cost Hamiltonian:
${f(\beta,\gamma)=\langle\psi(\beta,\gamma)|H_{C}|\psi(\beta,\gamma)\rangle.}$ (2)
As l increases ($l\to\infty$), QAOA can theoretically converge to the optimal solution. Increasing $l$ generally improves the solution quality but also increases the complexity of parameter optimization and susceptibility to hardware errors. Hence, in practice, $l$ is limited by the trade-off between the achieved approximation quality and the computational or hardware constraints [45, 17].
@@ 4.2 QUBO - Representing Optimization Problems
$\bar{\textbf{{x}}}={\rm arg~{}\min_{\textit{x}}}~{}\bar{y},$ (3)
$\bar{y}=\sum_{\textbf{{x}}_{i}\in\{0,1\}^{n}}\textbf{{x}}^{T}\mathbf{Q}\textbf{{x}}.$ (4)
@@ 4.3 Building QUBOs
$y=w_{0}+\sum_{i=1}^{n}w_{i}{\textit{x}}_{i}+\frac{1}{2}\sum_{f=1}^{m}\left[\left(\sum_{i=1}^{n}v_{i,f}{\textit{x}}_{i}\right)^{2}-\sum_{i=1}^{n}v_{i,f}^{2}{\textit{x}}_{i}^{2}\right].$ (5)
@@
To analyze the performance of QAOA, dq-QAOA, and DQAOA, we generate QUBO problems after training FM with datasets that depict the relationship between material structures (spectral filters) and their calculated FOMs. Hence, these QUBO matrices are large and dense (Fig. S5), reflecting real-world optimization scenarios. In this study, the QUBO size (i.e., the length of a binary bit string representing a material structure) ranges from 6 to 1,000, indicating 6-bit to 1,000-bit problems.
@@ 4.4 QUBO Decomposition
Large QUBO problems are hard to solve with current conventional quantum technologies. Therefore, we decompose a large QUBO into smaller sub-QUBOs and aggregate sub-solutions from the sub-QUBOs to obtain a global solution. This decomposition strategy greatly reduces the complexity of handling a large QUBO. For example, a QUBO problem with 8 decision variables involves 8 self-interactions and 28 pairwise interactions. By fixing 3 decisions and decomposing the QUBO into several sub-QUBOs with 5 decisions each, we can reduce the complexity of each sub-QUBO to have 5 self-interactions and 10 interactions (Fig. 1d), making it more suitable for current quantum computing systems. This QUBO decomposition can find a true global optimal solution of an original large QUBO when preserving essential structure and interactions of the original problem [47, 48]. However, it can be computationally expensive and ineffective to solve all sub-QUBOs to identify the global optimal solution, especially for large-scale problems due to the exponential number of possible combinations ${}_{\textbf{{n}}}C_{\textbf{{k}}}$. Here, n and k represent the size of the original QUBO and sub-QUBOs, respectively. Hence, our algorithm iterates the decomposition and aggregate processes with randomly selected sub-QUBOs among the whole possibilities. Then, sub-QUBOs are solved using DQAOA.
Divide-and-conquer methods with QAOA (DC-QAOA) have been proposed to tackle large-scale problems by partitioning them into smaller sub-problems. However, these strategies are primarily designed for sparse problems, such as Max-Cut, with the lack of considering correlated interactions between sub-problems. As a result, they are not effective for problems with densely (fully) connected interactions. In contrast, the decomposition and aggregation policies implemented in DQAOA demonstrate significantly better performance for real-world problems characterized by dense and large configurations, as shown in Figs. 4a,b and S3 [26, 27, 15]. For a comprehensive comparative analysis of approximation ratios and time-to-solution across different methods, we formulate five distinct QUBOs with random elements following a Gaussian distribution (mean = 0, standard deviation = 0.15). These QUBOs exhibit similar element distributions to those obtained from the FM training with material data, effectively representing the characteristics of real-world optimization problems (Fig. 4a,b, and Fig. S5). For DC-QAOA, the sub-problem size that yields the highest approximation ratio is selected from the range of 4 to 22, leading to variations in time-to-solution based on the chosen sub-problem size.
4.5 HPC Systems for Computational Experiments
@@
Our algorithm is tested on the Frontier supercomputer at the Oak Ridge Leadership Computing Facility, which is equipped with 64-core AMD “Optimized 3rd Gen EPYC” CPUs and 512 GB of memory per compute node [49]. In addition, each core interfaces with quantum systems (specifically, Qiskit Aer simulator in this study). dq-QAOA runs with a single core while DQAOA exploits multiple cores and nodes depending on the problem size. In this work, we use 50 cores per node, meaning that 150-bit problems would require 150 cores across 3 compute nodes. Distributed tasks across multi-cores and nodes are achieved by a message passing interface (MPI) implementation using mpi4py, enabling the simultaneous handling of numerous sub-QUBOs in each iteration [50].
For the implementation of dq-QAOA and DQAOA on the QCSC architecture, we utilize the Defiant system, which provides the advantages of extended wall times and reduced queuing times. The Defiant system features a configuration similar to Frontier, equipped with 64-core AMD EPYC 7662 “Rome” CPUs and 256 GB of memory per compute node.
4.6 Quantum Hardware
To use quantum hardware, we utilize a new version of qiskit (1.2.4) and qiskit-ibm-runtime (0.32.0) to interface with IBM quantum devices as backends. The QAOA ansatz is constructed using the QAOAAnsatz function and subsequently transpiled using a generate-preset-pass-manager provided by qiskit. The number of QAOA layer is set to 1 and the optimization level for circuit transpilation is set to 2.
We use FakeBrisbane and FakeKyiv as emulated backends (fake backends), which simulate the behavior of a quantum device while incorporating the effects of hardware errors due to noise. For quantum hardware, we utilize the devices IBM-Strasbourg and IBM-Kyiv, each featuring 127 qubits. Among these devices, the least busy device is selected for quantum circuit execution. The time-to-solution is calculated as the sum of quantum processing unit (QPU) usage time and parameter optimization time on classical hardware, explicitly excluding queueing delays. The time-to-solution on the simulators and emulators scales exponentially with problem size. In contrast, the time-to-solution of quantum hardware exhibits linear scaling, highlighting the potential quantum advantage over quantum simulation for solving large-scale problems (Fig. 1c).
We demonstrate the performance of dq-QAOA and DQAOA on IBM-Strasbourg and IBM-Kyiv, which provide a direct interface between HPC cores/nodes and quantum computing systems, enabling the distribution of sub-QAOA tasks across a quantum-centric supercomputing architecture (Fig. 4c). Hardware noise, particularly affecting two-qubit gates, makes solving QUBOs with n $>$ 22 on IBM quantum devices challenging, often resulting in poor performance, including negative approximation ratios. Emulators are similarly constrained, being difficult to handle QUBOs with n $>$ 12 due to exponential memory demands inherent in density matrix simulation. To address these constraints, we solve a QUBO with n = 22 with a sub-QUBO size of 11 as a proof-of-concept. Given the limited quantum hardware access, the number of iterations is restricted to 5. Directly solving the QUBO with QAOA on the quantum device results in a low approximation ratio of 0.6740 due to hardware noise and deep circuit depth. However, the use of DQAOA and dq-QAOA significantly improves the approximation ratios to 0.9727 and 0.9365, respectively (Fig. 4d,e). These findings highlight the practical feasibility of our distributed approach for addressing large and dense optimization problems on quantum hardware (Fig. 4d). For large-scale testing (n = 100), we omit the initialization step in both dq-QAOA and DQAOA to manage quantum computing costs. Instead, the global solution is initialized with random binary variables (Fig. 4f), and we extend the number of iterations up to 40. This approach can result in high approximation ratios.